OpenStudy (anonymous):

find the slope of the line that contaiins the points(-4,-3) and (5,6)

OpenStudy (anonymous):

1

OpenStudy (anonymous):

did you use the slope formula?

OpenStudy (anonymous):

\[slope = \frac{\Delta y}{\Delta x} = \frac{y_2-y_1}{x_2-x_1} = \frac{6+3}{5+4} = 1\]

OpenStudy (anonymous):

m=y2-y1 __________ x2-x1

OpenStudy (anonymous):

http://cs.selu.edu/~rbyrd/math/slope/